A 30ft ladder weighing 100 lb having its center of the mass one-third of the way up from the bottom rests against a smooth wall so that it makes an angle of  with the ground. If the coefficient of friction between the ground and the ladder is 0.4, how high can a 150-lb man go before the ladder slips?



1
Expert's answer
2022-04-11T12:02:41-0400

"mg\frac l3 \cos \alpha +Mgx\cos\alpha=N_l l\sin\alpha,"

"g\cos\alpha (Mx+m\frac l3)=N_l l\sin \alpha,"

"Mx+m\frac l3=\frac{N_l l\tan\alpha}{g} \leqslant \\ \mu(m+M)l \tan\alpha,"

"x\leqslant l(\mu\ tan \alpha(1+\frac mM)-\frac{m}{3M})=28~\text {ft}."